OpenStudy (anonymous):

solve the equation \(2x^4+3x^3-4x^2-3x+2=0\)

OpenStudy (anonymous):

first i what i did is below 2\((x^4+1\))+3\((x^3-x\))-\(4x^2\)=0

OpenStudy (anonymous):

multiplying both sides by \(\dfrac1x^2\) 2\((x^2+\dfrac{1}{x^2}\))+3\((x-\dfrac{1}{x}\))-4=0 Next what we will do

OpenStudy (anonymous):

@phi @math&ing001 @Mokeira @midhun.madhu1987 @fabiomartins @Abhisar @hero @helper99

OpenStudy (mokeira):

http://www.geteasysolution.com/2x%5E4-3x%5E2-2=0

OpenStudy (anonymous):

i did google ..but i don't understand @Mokeira

OpenStudy (anonymous):

@midhun.madhu1987 @sasogeek

OpenStudy (mokeira):

neither do I...let me try it though and see if i can help you

OpenStudy (anonymous):

@yayanieves @Jaynator495 @StarBreakerMegg @cj49

OpenStudy (anonymous):

@MilenaSaeger

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

im sorry but I cant help

terenzreignz (terenzreignz):

Try Rational Root test.

OpenStudy (anonymous):

@bibby @undeadknight26

OpenStudy (asnaseer):

I usually try to spot patterns in questions like this. For example in your equation:\[2x^4+3x^3-4x^2-3x+2=0\]I would first try to factor the \(x^4\) and \(x^2\) terms together and then the \(x^3\) and \(x\) terms to get:\[2x^2(x^2-2)+3x(x^2-1)+2=0\]Now I can see that I should try and make both the first 2 terms have \((x^2-1)\) so that I can factor it out. So I woud proceed as follows:\[2x^2(x^2-1)-2x^2+3x(x^2-1)+2=0\]Hopefully you can continue from here...

OpenStudy (anonymous):

no @asnaseer i am dong not in this way not by factorization by taking other way

OpenStudy (asnaseer):

@ziqbal103 - please explain what you mean - I don't quite understand what you are trying to say?

terenzreignz (terenzreignz):

Ohh... right. dangit, asnaseer... I really hate finding roots to long polynomials XD

OpenStudy (cj49):

x=-1,-2,1

OpenStudy (cj49):

https://www.youtube.com/watch?v=xY7Vpxvm62A

OpenStudy (anonymous):

great link

OpenStudy (anonymous):

@triciaal @taylor12344 @Luigi0210 @liliy

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

did you check the youtube link cj posted?

OpenStudy (anonymous):

yes but i prepare these uestion not in this way that @cj49 sent the link...@zimmah

OpenStudy (anonymous):

well, i wouldn't know how to do it any other way

OpenStudy (anonymous):

except calculator

OpenStudy (anonymous):

i am going to copying (\(x+\dfrac{1}{x}\))=t , then \((x^2+\dfrac{1}{x^2}-2\))=\(t^2\) or \(x^2+\dfrac{1}{x^2}\)=\(t^2+2\)

OpenStudy (anonymous):

in this method i want to solve my question

OpenStudy (asnaseer):

in that case you would be better off using:\[t=x-\frac{1}{x}\]

OpenStudy (anonymous):

yes do u know this method @asnaseer

OpenStudy (asnaseer):

which also gives:\[t^2+2=x^2+\frac{1}{x^2}\]

OpenStudy (asnaseer):

@ziqbal103 - no I have never seen this method. Does it have a name?

OpenStudy (anonymous):

equation in exponenetial form @asnaseer

OpenStudy (asnaseer):

if you use the substitutions I gave above, your equation reduces to:\[2(t^2+2)+3t-4=0\]which is a simple quadratic.

OpenStudy (anonymous):

yes but in above i don't know y they take square root of \(x-\dfrac1{x}\)

OpenStudy (asnaseer):

what do you mean?

OpenStudy (triciaal):

2x4+3x3−4x2−3x+2=0 I don't have time to look at all the responses one approach is to apply the remainder theorem to start and then factorize if the value of x is a factor then the remainder = 0 2 * 2 = 4 factors of 4 1, -1. 2, -2 f(1) = 2 + 3 - 4 - 3 +2 = 0 ; x=1 (x-1) is a factor f(-1)= 2 - 3 -4 + 3 + 2 = 0; x = -1 (x + 1) is a factor f(2) = f(-2)= the degree of polynomial and the rule of signs will let you know how many roots you can have in this case maximum of 4

OpenStudy (asnaseer):

@ziqbal103 - you ended up with:\[2(x^2+\frac{1}{x^2})+3(x-\frac{1}{x})-4=0\]if we use the substitution:\[t=x-\frac{1}{x}\implies t^2=x^2+\frac{1}{x^2}-2\]we get:\[2(t^2+2)+3t-4=0\]\[\therefore 2t^2+4+3t-4=0\]\[\therefore 2t^2+3t=0\]\[\therefore t(2t+3)=0\]\[\therefore t=0\text{ or }-\frac{3}{2}\]so next you would use:\[t=x-\frac{1}{x}\]to solve for \(x\) for each value of \(t\)

OpenStudy (triciaal):

repost of @asnaseer which also gives: t2+2=x2+1x2 solve this quadratic will give t = 0 , t = -3/2 x = 1, -1, 1/2, -2

OpenStudy (anonymous):

in this question 6x^4-35x^3+62x^2-35x+6 in this multiplying by both sides by \(\dfrac1{x^2}\) we get 6\((x^2+\dfrac1{x^2}\))-35\((x-\dfrac1{x}\))+62=0 let \(x-\dfrac1{x}\)=t ,then \(x^2+\dfrac1{x^2}+2\)=\(t^2\) in above why we don't take the square of \(x-\dfrac1{x}\)

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

@goformit100 please halp me

OpenStudy (asnaseer):

\[(x-\frac{1}{x})^2=x^2+\frac{1}{x^2}-2\]so we ARE squaring it

OpenStudy (anonymous):

but in my guide the procedure is diff. and i don't understand..i solved it but now i forget @asnaseer

OpenStudy (asnaseer):

I am finding it very difficult to understand where you are getting stuck. What is this "guide" that you are following? Is it some text book?

OpenStudy (anonymous):

yes it is text book @asnaseer

OpenStudy (asnaseer):

It might help if you uploaded a picture of the page describing the procedure that you are following

OpenStudy (anonymous):

it's a pretty complicated method they use in that textbook

OpenStudy (asnaseer):

what is the name of this text book?